OpenStudy (anonymous):

factoring.. x^4+2x^2+1

hero (hero):

Hint: If you let x^2 = u, then you will have u^2 + 2u + 1

OpenStudy (anonymous):

ok, i got that part. now it asks for me to find the zeros and their multiplicity. i know the zeros are i and -i, how do i know their multiplicity?

OpenStudy (anonymous):

@hero

hero (hero):

u^2 + u + u + 1 u(u + 1) + 1(u+1) (u+1)(u+1) (u+1)^2 Replace u with x^2 (x^2 + 1)^2 Set (x^2 + 1)^2 = 0 Take square root of both sides: x^2 + 1 = 0 x^2 = -1 x = + or - sqrt{-1} x = + or - i Okay